A delivery truck leaves a town traveling at 35 miles/hour. An hour later an emergency vehicle sets off after the truck from the same place. It takes the same route as the truck but travels at 55 miles/hour. How long will it take the emergency vehicle to catch up with the truck?

Answer by richwmiller(17219)

the truck travels 35 miles in one hour
the emergency vehicle is traveling 20 mph faster
20*t=35
t=35/20=7/4= 1.75 hours